consider the roll of a pair of dice, and let X denote the sum of the two numbers appearing on the dice. find the probability distribution of X and find the cumulative distribution function. what is the most likely sum
1
Expert's answer
2019-11-07T09:33:24-0500

The first step is calculating the sample space S= {2, 3, ,3 ,4,4,4 ,5,5,5,5,6,6,6,6,6,7,7,7,7,7,7,8,8,8,8,8,9,9,9,9,10,10,10,11,11,12}

count the frequency of each possible outcome and divide with total count n =36 to find PDF then calculate CDF for each X by adding the current and all preceding PDFs.


X PDF CDF

2 "\frac 1 {36}" "\frac 1 {36}"


3 "\frac 1 {18}" "\frac {1} {12}"


4 "\frac 1 {12}" "\frac 16"


5 "\frac 19" "\frac 5 {18}"


6 "\frac 5 {36}" "\frac 5 {12}"


7 "\frac 1 6" "\frac 7 {12}"


8 "\frac 5 {36}" "\frac {13} {18}"


9 "\frac 1 9" "\frac 5 6"


10 "\frac 1 {12}" "\frac {11} {12}"


11 "\frac 1 {18}" "\frac {35} {36}"


12 "\frac 1 {36}" 1


The most likely sum is 7 it has the highest PDF .